Compute the discounted payback statistic for Project C if the appropriate cost of capital is 8 percent and the maximum allowable discounted payback period is three years. (Do not round intermediate calculations and round your final answer to 2 decimal places.) Project C

Time: 0 1 2 3 4 5

Cash flow: −$2,200 $960 $840 $880 $540 $340
Answer #1

Discounted payback period is the time in which initial investment comes back as cash inflow.

Discounted Cash Flow in Year 1 = 960/(1.08) = $888.89

Discounted Cash Flow in Year 2 = 840/(1.08)2 = $805.90

Discounted Cash Flow in Year 3 = 880/(1.08)3 = $698.57

Discounted Cash Flow in Year 4 = 540/(1.08)4 = $396.92

Discounted Payback period = 1(888.89) + 1(805.90) + 0.72(505.21/698.57)

Discounted Payback Period = 2.72 years